A car can be brought to rest from a speed of 200 m/s in a time of 2 seconds.Calculate the deceleration



1
Expert's answer
2021-05-05T10:54:25-0400

By definition, the acceleration (negative deceleration) is given as follows:


"a = \dfrac{v_f-v_i}{t}"

where "v_f = 0m\/s" is the final speed of the car, "v_i = 200m\/s" is its initial speed, "t = 2s" is the time passed. Thus, obtain:


"a = \dfrac{0-200m/s}{2s} = -1m/s^2"

Negative acceleration is called deceleration. Thus, the deceleration is "1m\/s^2".


Answer. "1m\/s^2".